Lesson 13 Concentration Problems

  1. What is the concentration, in g/mL, of a solution made by dissolving 75.0 g of phosphoric acid in 50.0 mL of solution?
  2. What mass of NaCl is needed to produce 500 mL of a 2.0 mol/L solution?
  3. What is the molarity of a solution that contains 0.350 g of barium fluoride, BaF2, in 0.50 L of solution?
  4. What is the molar concentration of a solution containing 10.0 g of NaOH in 250 mL of solution?
  5. What mass of silver nitrate, AgNO3, is needed to make 0.50 L of a 0.10 mol/L solution?
  6. What volume of 1.0 mol/L solution of CuS can be made from 23.9 g of CuS?
  7. A 100 mL sample of a 6.0 mol/L solution of H2SO4 is diluted to a final volume of 500 mL. What is the concentration of the diluted solution?
  8. What volume of a 6.0 M NaOH solution is needed to produce 3.0 L of a 0.1 M solution?
  9. Explain the steps you would follow to make 500 mL of a 0.5 mol/L solution of sodium chloride (NaCl) in the lab.
1
Expert's answer
2021-03-24T01:07:35-0400

1. 1.5 g/ml

2. 58.5 g

3. 0.004 M

4. 1M

5. 8.5 g

6. 250 ml

7. 1.2 M

8. 50 ml

9. n(NaCl) = Cm×V

n(NaCl) = 0.5×0.5 = 0.25 mole

m(NaCl) = n×M

m(NaCl) = 0.25×58.5 = 14.6 g

Making:

1. Weight 14.6 g of NaCl and put it into the glass.

2. Dilute it with 400 mL of water.

3. Decant solution in the graduated cylinder or other glassware and put rests of water.

4. Final volume should be 500mL.